Straight bank loan. Right Bank offers EAR loans of 8.81 % and requires a monthly payment on all loans. What is the APR for these monthly​ loans? What is the monthly payment for a loan of ​(a​) $215,000 for 6 ​years, ​(b​) $435,000 for 11 ​years, or ​(c​) $ 1,200,000 for 29 ​years?

What is the APR for these monthly​ loans? _______% ​(Round to three decimal​ places.)

​(a​) What is the monthly payment if a loan is for $215,000 for 6 ​years? ​$ _________ (Round to the nearest​ cent.)

​(b​) What is the monthly payment if a loan is for $435,000 for 11 ​years? ​$ ________ ​(Round to the nearest​ cent.)

​(c​) What is the monthly payment if a loan is for $1,200,000 for 29 ​years? ​$ __________ ​(Round to the nearest​ cent.)
Answer #1

Answer:

EAR=8.81%

APR=12*((1+EAR)^(1/12)-1)

APR=12*((1+8.81%)^(1/12)-1)

APR=8.47%

A)

Loan amount L=$215000

Tenure = 6 years

N=6*12=72 months

r=APR/12=0.71%

So per month payment = L*r/(1-(1+r)^-N)=215000*0.71%/(1-(1+0.71%)^-72)=$3819.50

B)

Loan amount L=$435000

Tenure = 11 years

N=11*12=132 months

r=APR/12=0.71%

So per month payment = L*r/(1-(1+r)^-N)=435000*0.71%/(1-(1+0.71%)^-132)=$5077.20

C)

Loan amount L=$1200000

Tenure = 29 years

N=29*12=348 months

r=APR/12=0.71%

So per month payment = L*r/(1-(1+r)^-N)=1200000*0.71%/(1-(1+0.71%)^-348)=$9274.57